StatChapter03 Part 2: Problem 7 50 66 116 33 51 84 07 134 23 You haw
Answer #1

a)

probability=116/231 =0.5022

b)

probability=97/231=0.4199

c)probability=66/231=0.2857

d)

probability=(116+134-66)/231=184/231=0.7965

e)

probability=14/231=0.0606